Consider the polynomial function 𝑃(π‘₯) = 2π‘₯^3 βˆ’ π‘šπ‘₯^2 + π‘₯ βˆ’ 5π‘š. The remainder when P(x) is divided by (π‘₯ βˆ’ 2) is four times the remainder from dividing P(x) by (π‘₯ + 1). Determine m algebraically and show all your work.Β 


1
Expert's answer
2021-11-26T14:51:38-0500

"Let\ P(x)=2x^3-mx^2+x-5m"

The remainder when P(x) is divided by (x-2) is;

"P(2)=2\cdot8-m\cdot4+2-5m"

"=16-4m+2-5m"

Thus: P(2)=18-9m

And the remainder when P(x) is divided by (x+1) is;

"P(-1)=-2-m-1-5m"

=-3-6m

P(-1)=-3-6m

From the question: P(2)=4*P(-1)

18-9m=4(-3-6m)

18-9m=-12-24m

-9m+24m=-12-18

15m=-30

"m=\frac{-30}{15}"

m=-2